Circular No. 7305

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999em IN NGC 1637 D. C. Leonard, A. V. Filippenko, and R. T. Chornock, University of

California at Berkeley, write: “Preliminary reduction of spectropolarimetry (ranges 400–510 and 600–730 nm), obtained on Nov. 5 UT with the Shane 3-m reflector at Lick Observatory, indicates very weak continuum polarization (p < 0.3 percent, uncorrected for interstellar polarization) with no significant polarization variation across the strong Balmer lines (δp < 0.1 percent). When modeled in terms of an oblate, electron-scattering atmosphere (cf. Hoflich 1991, A.Ap. 246, 481), this suggests a nearly spherical continuum-scattering region, or else a nearly pole-on viewing orientation.”

COMET C/1999 V1 (CATALINA) Additional astrometry appears on MPEC 1999-V29, from whence the

following preliminary orbital elements are taken:

T = 1999 Sept.15.442 TT  ω = 176.^{o}561
Ω = 295.056 }2000.0
q = 3.06110 AU  i = 15.206
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Nov. 8 9^{h}39.^{m}60 +11^{o}07.^{′}5 3.091 3.106 81.^{o}7 18.^{o}4 17.4
13 9 44.04 +10 19.8 3.034 3.115  85.4 18.5 17.3
18 9 48.07 + 9 32.9 2.977 3.124  89.2 18.4 17.3
23 9 51.70 + 8 46.9 2.921 3.134  93.2 18.3 17.3
28 9 54.91 + 8 02.0 2.865 3.145  97.2 18.1 17.3
Dec. 3 9 57.66 + 7 18.4 2.811 3.157 101.4 17.8 17.2
8 9 59.94 + 6 36.2 2.758 3.169 105.6 17.4 17.2
13 10 01.74 + 5 55.8 2.707 3.182 110.0 16.9 17.2
18 10 03.05 + 5 17.3 2.659 3.195 114.5 16.3 17.2
23 10 03.86 + 4 40.9 2.614 3.210 119.1 15.5 17.2
28 10 04.16 + 4 06.8 2.571 3.225 123.9 14.7 17.1

COMET 141P/MACHHOLZ 2 Total visual magnitude estimates for component D: Oct. 31.74 UT,

13.8 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); Nov. 7.24, 12.3 (N. Biver, Oahu, HI, 0.26-m reflector); 8.22, 12.5 (Biver); 9.09, 11.9: (A. Hale, Cloudcroft, NM, 0.41-m reflector).

1999 November 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green